OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x^3y^4)^2

OpenStudy (anonymous):

\[(3x^{3} * y^{4})^2\] is that what it looks like...?

OpenStudy (anonymous):

Part one: Three times a number cubed times a number to the fourth power all squared

OpenStudy (anonymous):

@gaara438125 Yes thats what it looks like!

OpenStudy (anonymous):

\[(3x^{3} * y^{4})(3x^{3} * y^{4})\] wanna FOIL that out? XD

OpenStudy (anonymous):

@kmalone99 can you help explain how u got those so i understand it better? Please.

OpenStudy (anonymous):

\[9x ^{6}y ^{8}\]

OpenStudy (anonymous):

@kmalone99 guide him to the answer not just give it to him http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

@kmalone99 that's not even right....

OpenStudy (anonymous):

Im on the last question for my review guide and the hours i can call my teachers right now you cant call past 8 so, Im looking for someone to help explain this to me.

OpenStudy (anonymous):

yes thats how it reads the number 3 times x which can be any number to the third power which means anumber times itself three times and times y a another number times itself four times

OpenStudy (anonymous):

gaara438125 whata not right?

OpenStudy (anonymous):

@kmalone99 the: 9x^6y^8

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

no -.-

OpenStudy (anonymous):

\[(3x^{3} * y^{4})(3x^{3} * y^{4})\] you do the First and Last what about outer and inner?

OpenStudy (anonymous):

what do you think it is old great one? :-)

OpenStudy (anonymous):

\[9x^6 + 6x^3y^4 + y^8\]

OpenStudy (anonymous):

Can you guys explain this to me, so i know how to do it?

OpenStudy (anonymous):

I know my answer is right, it does not say plus in the middle so check again. so just call me Mster Malone

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[3x^3 * y^4 = 3x^3y^4\]

OpenStudy (anonymous):

(3x^{3} * y^{4})(3x^{3} * y^{4}) do you know how to foil that....? like do you honestly know?

OpenStudy (anonymous):

\[(3x^{3} * y^{4})(3x^{3} * y^{4})\]

OpenStudy (anonymous):

lets do this what do you not understand on what I wrote? answer gaara is not correct

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (anonymous):

okay for this, xplain, using complete sentences, how to simplify the expression below. Dont you have to write out explainging how you simplify it? not just writing it in word form?..

OpenStudy (anonymous):

\[(3x^{3} * y^{4})^2 = (3x^{3} * y^{4})(3x^{3} * y^{4})\] then you FOIL

OpenStudy (anonymous):

sorry gaara but your answer was wrong

OpenStudy (anonymous):

wtf is FOIL, i was never taught FOIL??

OpenStudy (anonymous):

@kmalone99 facepalm at your right now...learn something and @letsdothis15 FOIL is when you have a binomial you multiply First, Outer, Inner, and Last

OpenStudy (anonymous):



OpenStudy (anonymous):

letsdothis it is not foil any way.....the question says to explain

OpenStudy (anonymous):

you multiply the lines to what they're connected to. the 1-4 are in which order

OpenStudy (anonymous):

@kmalone99 "Part 2: What is the simplified expression?"

OpenStudy (anonymous):

omg im so confused. I got all the other problems on my lesson just not this. -.-

OpenStudy (anonymous):

for part 2 \[(3x^{3} * y^{4})^2\] that also looks like \[(3x^{3} * y^{4})(3x^{3} * y^{4})\]

OpenStudy (anonymous):

gaara your not right 3x3=9 and x^3 times x^3 is=x^6 and y^4 times y^4 =y^8

OpenStudy (anonymous):

@kmalone99 you realize (3x^{3} * y^{4}) is being squared....?

OpenStudy (anonymous):

letsdothis I was a high school math teacher for 5 years I have a Bachlors degree in Math and I'm in Graduate school working on my Masters in Math.

OpenStudy (anonymous):

yet you can't understand the simple concept of multiply binomials....cool

OpenStudy (anonymous):

gaara you are wrong please go check how to multiply variables with expont.

OpenStudy (anonymous):

it's binomials....

OpenStudy (anonymous):

The answer i got is what kmalone99. so you multiply the 3&3 which is 9. then you add the exponent 3&3 also then u get 6. then u add the 4&4 which is 8?

OpenStudy (anonymous):

The lesson I am in is,called, Multiplying Monomials.